Susan is a saleswoman. She is paid a base salary of $1600 a month plus a 4% commission on the sales she makes. What was her sales amount last month if she earned a total of $2487.84 ?






Answer by Mathpassionate(25)

Susan is a saleswoman. She is paid a base salary of $1600 a month plus a 4% commission on the sales she makes. What was her sales amount last month if she earned a total of $2487.84 ?


Let A = Sales Amount
Basic income (B) = 1600
Rate (R) = 4%
Commission (C) = 4*A/100
Total monthly income (T) = 2487.84



T = B + C
T = B + 4A/100
2487.84 = 1600 + 4A/100
2487.84 - 1600 = 4A/100
887.84 = 4A/100
(887.84)/4 = A/100
A= (221.96)*100
A= 22196

So, the sales amount was $22.196 which is the answer.

If you have any questions, please let me know.